The PI's presented the status of their instruments (see annex 6).
Investigating degradation/calibration issues after the SOHO interruption last year.
Discussed the remaining life time of detector A and B. With continuous use they estimate the detector A and B to last for about 2 and 3 months, respectively. A fallback mode is being investigated. It will utilize the lower 30 pixels of the detector, which has not yet been exposed to light. The upper unexposed part can not be reached.
A drift in solar images was reported from the SUMER team. It was in the range 2 arcseconds in solarX and 0.5 arcsec in solarY. This is not consistent with MDI for the same period (approximately 1 arcsec in X).
CDS is almost fully booked until November. GIS detector 2 is not in use at the moment. Very little scientific impact since NIS is used by most people. Have plans for the upcoming eclipse, the Mercury transit and the Whole Sun Month.
CDS has been mentioned as a source of pointing disturbances for other instruments. Instrument teams should look into the period 9-14 June 1999. In this period CDS did not do any OPS movements at all so this could be a nice check.
No effects on the CCD was found after 4 hours of shutterless operation.
Presented results from the press release (detection of active regions on far side of the Sun + shadow of comet Hale-Bopp). A question was raised about when one can predict active regions on the far side of the Sun on a regular bases. This is a manpower problem but CNES has indicated interest to support such a study.
Noted that there were no SOHO key parameters on the CD-rom from ISTP since October 1998. L. Sanchez will investigate this.
Test on C1 still ongoing